EDITORIAL

Terry Goodman

We want to acknowledge Robert E. Kennedy who has served as the Coordinating Editor for the MJMS since its inception. Bob has had the large task of distributing manuscripts to referees and coordinating and scheduling the publication of the accepted articles. Bob has just entered into retirement and we wish him all the best. We appreciate his consistently positive contribution to the MJMS.

We are pleased that Shing So will assume the role of Coordinating Editor and we know that he will continue the high quality of work that Bob has given the Journal. Also, we want to welcome two new associate editors: Phoebe Ho McLaughlin from Central Missouri State University as the Number Theory and Algebra editor and B. Dale Daniel from Lamar University as the Topology and Geometry editor.

Once again we want to use our editorial space for this issue to thank and acknowledge those individuals who serve as referees for the MJMS. We believe one of the strengths of our journal is the use of a thorough and consistent review process. The persons listed below play a vital role in that review process and we appreciate and applaud their efforts. any reader who is interested in serving as a referee for the MJMS should complete the Referee Applicaton Form found after the Announcements.

List of Referees

Hadi Y. Alkahby, Dillard University Jamuna P. Ambasht, Benedict College Mohammad K. Azarian, University of Evansville Seung–Jin Bang, Seoul, Korea Gerald Bergum, South Dakota State University John E. Bernard, University of Texas–Pan American Jim Bruening, Southeast Missouri State University Joseph Chance, University of Texas–Pan American Hang Chen, Central Missouri State University Eungchun Cho, Kentucky State University Gregory Davis, University of Wisconsin–Green Bay Bryan Dawson, Union University Richard Delaware, University of Missouri–Kansas City Joseph B. Dence, University of Missouri–St. Louis Albert Dixon, College of the Ozarks

List of Referees (cont.)

Clayton Dodge, University of Maine John J. Edgell, Jr., Southwest Texas State University Vince Edmondson, Central Missouri State University Larry Eifler, University of Missouri–Kansas City Ekemezie J. Emeka, Columbia College Martin Erickson, Truman State University Russell Euler, Northwest Missouri State University Joe Flowers, Trinity University Jayanthi Ganapathy, University of Wisconsin–Oshkosh Ramesh Garimella, Northwest Missouri State University Raymond J. Grinnell, University of the West Indies Victor Gummersheimer, Southeast Missouri State University K.L.D. Gunawardena, University of Wisconsin–Oshkosh Dennis R. Harmon, Missouri Southern State College Stan Hartzler, Virginia Polytech Institute & State University Troy Hicks, University of Missouri–Rolla Joe Howard, New Mexico Highlands University Timothy P. Keller, George Mason University N. J. Kuenzi, University of Wisconsin–Oshkosh Charles E. Lamb, University of Texas–Austin Mulatu Lemma, Savannah State University Man Chun Leung, National University of Singapore Yi–Hsin Liu, University of Nebraska–Omaha William Livingston, Missouri Southern State College Roger Marty, Cleveland State University Wayne McDaniel, University of Missouri–St. Louis Gary McDonald, Northwest Missouri State University Rhonda L. McKee, Central Missouri State University Phoebe McLaughlin, Central Missouri State University Peter R. Mercer, St. Mary’s College of Maryland Kandasamy Muthvel, University of Wisconsin–Oshkosh Ollie Nanyes, Bradley University Krzysztof Ostaszewski, University of Louisville Miguel Paredes, University of Texas–Pan American Don Redmond, Southern Illinois University Jeffrey A. Rosoff, Gustavus Adolphus College Prasanna Sahoo, University of Louisville Jau–Shyong Shiue, University of Nevada–Las Vegas

List of Referees (cont.)

Donald P. Skow, University of Texas–Pan American Randall J. Swift, Western Kentucky University Jinhua Tao, Central Missouri State University Songlin Tian, Central Missouri State University Mohan Tikoo, Southeast Missouri State University Mathew Timm, Bradley University Alvin Tinsley, Central Missouri State University Peter L. Vachuska, University of Wisconsin–Washington Co. Joseph Wiener, University of Texas–Pan American Bobby Neal Winters, Pittsburg State University Fred Worth, Henderson State University